OpenStudy (anonymous):

What is the volume of a sphere with radius 5?

OpenStudy (anonymous):

A. 33.33 units3 B. 126.67 units3 C. 100 units3 D. 166.67 units3

OpenStudy (john_es):

\[V=\frac{4}{3}\pi R^3\] Use R=5, and calculate.

OpenStudy (anonymous):

\[V=4/3\pi25\]

OpenStudy (anonymous):

am i going good lol

OpenStudy (john_es):

Not exactly, is R^3, so R^3=125

OpenStudy (anonymous):

4/3pi125

OpenStudy (anonymous):

?

OpenStudy (anonymous):

what i do from here ?

OpenStudy (john_es):

Ok, then now, calculate, \[4/3*3.14*125\approx523.6\] This is the correct result, so I don't know why you give the options before.

OpenStudy (anonymous):

i only have those options thats where im confused

OpenStudy (john_es):

But those options are incorrect. Do you have an image of the problem?

OpenStudy (anonymous):

this is all i got

OpenStudy (anonymous):

im still trying to figure it out im even looking on google too

OpenStudy (john_es):

Ok, now see, you left the pi out in your options. That was the problem. So you have to answer in units of pi. \[4/3∗\pi∗125≈166.67\pi\] So it is option D.

OpenStudy (anonymous):

you are correct ty ty